How would you solve...
Write the equation y = 1/4x-2 in standard form using only integers and a positive
coefficient for x.

Answer by Alan3354(69443)

Write the equation y = 1/4x-2 in standard form using only integers and a positive
coefficient for x.
--------------
Multiply by 4 to eliminate the fraction.
4y = x - 8
x = 4y + 8
x - 4y = 8